OpenStudy (anonymous):

Identify the axis of symmetry and the vertex of the fuction y=-2x^2+1

jimthompson5910 (jim_thompson5910):

In general, if we have y = ax^2 + bx + c the axis of symmetry is x = -b/(2a)

jimthompson5910 (jim_thompson5910):

In the case of y=-2x^2+1, a = -2, b = 0 and c = 1

OpenStudy (anonymous):

thanks i only put that ot do something more?

jimthompson5910 (jim_thompson5910):

what did you get for the axis of symmetry?

OpenStudy (anonymous):

i haven't done it

OpenStudy (anonymous):

but don't worry i al

OpenStudy (anonymous):

already got you thanks

jimthompson5910 (jim_thompson5910):

alright, yw